# ExoPredicator: Learning Abstract Models of Dynamic Worlds for Robot Planning

### v1

@@ Abstract
Long‑horizon embodied planning is challenging because the world does not only change through an agent’s actions: exogenous processes (e.g., water heating, dominoes cascading) unfold concurrently with the agent’s actions. We propose a framework for abstract world models that jointly learns (i) symbolic state representations and (ii) causal processes for both endogenous actions and exogenous mechanisms. Each causal process models the time course of a stochastic causal-effect relation. We learn these world models from limited data via variational Bayesian inference combined with LLM proposals. Across five simulated tabletop robotics environments, the learned models enable fast planning that generalizes to held‑out tasks with more objects and more complex goals, outperforming a range of baselines. [^1] Code: https://github.com/BasisResearch/predicators/releases/tag/ExoPredicator
@@ 1 Introduction
The standard planning representation used for decades, PDDL, also fails to capture this: it only models the effects of one’s own actions, and treats each action as instantaneous. Learning a symbolic PDDL planning model (Silver et al., 2023; Liang et al., 2024) combinatorially explodes with respect to temporal granularity. Vision-language models (VLMs) and vision-language-action models (Team et al., 2023; Black et al., 2025) could in principle reason about external causal mechanisms, but generalize poorly to novel situations, particularly when reasoning about temporal physical constraints.
@@ 3 Abstracting States, Time, and Causal Processes
We write $\mathcal{S}$ for the set of possible abstract states.
Causal Processes: Informal intuition. A causal process coarsely models a cause-effect relation. For example, opening a faucet above a bucket triggers a chain of cause-effect relations: a stopper opens, hidden pipes fill with water, water rises in the bucket, and eventually the bucket is filled. For decision making, many such details are irrelevant, or omitted from the state abstraction. A causal process therefore abstracts away such details by saying that certain conditions (the “cause”) later lead to other conditions (the “effect”). We further distinguish two kinds of causal processes. Endogenous processes correspond to the agent’s high-level actions or skills. They represent operations that are under the agent’s direct control, such as switching a faucet on. Exogenous processes describe the background dynamics of the environment. They represent events that are triggered by certain conditions but unfold without the agent’s continuous intervention, such as a kettle filling with water after being placed under a running tap. This separation allows the agent to reason about the consequences of its actions while also anticipating changes initiated by the environment itself.
@@
Abstract world models are equipped with a set of causal processes $\mathcal{L}$. Each causal process $L\in\mathcal{L}$ is defined by a schema tuple, $\langle\textsc{Par},C,O,E,W,p^{\text{delay}}\rangle$:
  • Par is a list of typed variables present in the condition or effect of the process (“parameters”).
@@
  • $W\in\mathbb{R}$ (“weight”) quantifies how likely the effect will happen when the conditions are satisfied.
  • $p^{\text{delay}}$ is a probability distribution over the delay between the process’s activation and effect.
Endogenous processes further include a skill $c\in\mathcal{C}$. A schema is instantiated into a ground causal process $\underline{L}=\langle\underline{\textsc{Par}},\underline{C},\underline{O},\underline{E},W,p^{\text{delay}}\rangle$ (and optionally $\underline{c}$) by substituting its parameters with specific objects. For example, the endogenous process SwitchBurnerOn (Figure 2) has parameters ?robot and ?burner. Its condition $C$ requires that the burner is off and the robot’s hand is free. Once executed, its effect $E$ (the burner being on) occurs after a delay sampled from $p^{\text{delay}}$.
Interdependence of state abstraction and causal processes. The right abstractions depend on downstream tasks, as our goal is generalization to unseen planning problems. Tasks demanding a more detailed state abstraction will generally require modeling more temporal dynamics. This couples the causal processes to the state abstraction.
Probabilistic Semantics. The causal processes define how abstract states evolve over time. Mathematically, $\mathcal{L}$ defines a probabilistic generative model over sequences of abstract states, $\left\{s_{t}\right\}$, where $t\in\mathbb{N}$ indexes fine-grained timesteps. We probabilistically model the effect of ground process $\underline{L}$ as a potential $E^{j}_{\underline{L}}$ upon each feature $j$ of the abstract state:
@@ 5 Process Learning and Predicate Invention
We initialize with 1-2 demonstration trajectories, and then perform online learning (Figure 3). Online learning involves planning to solve training tasks to collect further trajectories (when planning fails, random actions are taken). At each stage of online learning, we have a dataset of state-action trajectories $\mathcal{D}=\{(x_{0},\underline{c}_{0},\ldots,\underline{c}_{T-1},x_{T})\}$. Given predicates $\Psi$, this generates a dataset of abstract state trajectories $\{(s_{0},\ldots,s_{T})\}$. Given this dataset, we learn predicates (Section 5.3), exogenous processes (section 5.2), and continuous parameters (Section 5.1)—described in a reverse order, because each component is used in the previous component during learning.
@@ 5.1 Parameter Learning via Variational Inference
Each causal process has continuous parameters for the delay distribution $p^{\text{delay}}_{L}$ and the probabilistic weight $W_{L}$, and we have a global parameter $W_{F}$ for the frame axiom. Fixing all discrete structure—including the predicates—we seek parameters maximizing the marginal likelihood $p(\mathcal{D})=p(s_{1:t}|s_{0})=\sum_{\left\{\Delta_{1:t}^{\underline{L}}\right\}}p(s_{1:t},\{\Delta_{1:t}^{\underline{L}}\}|s_{0})$. This is intractable because there are combinatorially many ways of timing the cause-effects.
We therefore approximate the marginal likelihood by introducing variational distributions corresponding to the time at which each cause realizes its effect: distribution $q_{t}^{\underline{L}}(A_{t}^{\underline{L}})$ encodes beliefs about $A^{\underline{L}}_{t}$, the “arrival time” of the cause coming from $\underline{L}$ due to a cause at time $t$. This can be viewed as a change of basis where $A^{\underline{L}}_{t}=\Delta_{t}^{\underline{L}}+t$. Section A.3 derives the following variational lower bound on the marginal likelihood, which we optimize using Adam (Kingma, 2014).
$\log p\left(s_{1:t}|s_{0}\right)\geq\mathbb{H}_{q}\left[\{A_{t}^{\underline{L}}\}\right]+\sum_{t=1}^{T}\sum_{\underline{L}}C_{\underline{L}}(s_{1:t})\sum_{A_{t}^{\underline{L}}=t+1}^{T}q_{t}^{\underline{L}}(A_{t}^{\underline{L}})\log\left(p_{\underline{L}}^{\text{delay}}(A_{t}^{\underline{L}}-t)\right)+$
$\sum_{t,j}\log\left(F(s_{t}^{j}|s_{t-1}^{j})\right)+\sum_{\underline{L}}\sum_{t^{\prime}=1}^{t-1}q_{t^{\prime}}^{\underline{L}}(A_{t^{\prime}}^{\underline{L}}=t)C_{\underline{L}}(s_{1:t^{\prime}})O_{\underline{L}}(s_{t^{\prime}+1:t-1})\log\left(E_{\underline{L}}^{j}(s_{t}^{j})\right)$-
$\sum_{t,j}\log\sum_{\hat{s}_{t}^{j}\in\mathcal{S}^{j}}F(\hat{s}_{t}^{j}|s_{t-1}^{j})\prod_{\underline{L}}\prod_{t^{\prime}=1}^{t-1}\left(q^{\underline{L}}_{t^{\prime}}(A_{t^{\prime}}^{\underline{L}}=t)E_{\underline{L}}^{j}(\hat{s}^{j}_{t})^{O_{\underline{L}}(s_{t^{\prime}+1:t-1})}+\left(1-q^{\underline{L}}_{t^{\prime}}(A_{t^{\prime}}^{\underline{L}}=t)\right)\right)^{C_{\underline{L}}(s_{1:t^{\prime}})}$
@@ 5.2 Learning Exogenous Processes: Bayesian Model Selection, LLM Guidance
Given a set of trajectories $\mathcal{D}$, we would ideally learn the causal processes $\mathcal{L}^{\star}$ maximizing
$\mathcal{L}^{\star}$ $=\operatorname*{arg\,max}_{\mathcal{L}}p(\mathcal{L}|\mathcal{D})=\operatorname*{arg\,max}_{\mathcal{L}}p(\mathcal{D}|\mathcal{L})p(\mathcal{L})$ (intractable)
where $p(\mathcal{L})$ is a minimum description length prior and $p(\mathcal{D}|\mathcal{L})$ is approximated by Section 5.1. As this optimization is intractable, we learn a separate process $L_{\mathcal{C}}$ for each cluster $\mathcal{C}$:
$\mathcal{L}^{\star}$ $=\bigcup_{\mathcal{C}\in\textsc{Cluster}(\mathcal{D})}\{L_{\mathcal{C}}\}\quad\text{where }L_{\mathcal{C}}=\operatorname*{arg\,max}_{L}p(\mathcal{C}|L)p(L)$ (still intractable)
@@
$\mathcal{L}^{\star}$ $=\bigcup_{\mathcal{C}\in\textsc{Cluster}(\mathcal{D})}\left\{L_{\mathcal{C}}\right\}\quad\text{ where }L_{C}=\operatorname*{arg\,max}_{L\in\textsc{Prompt}(\mathcal{C})}p(\mathcal{C}|L)p(L)$ (tractable)
@@ 5.3 Learning State Abstractions: Program Synthesis and Local Search
$\Psi^{\star}$ $=\operatorname*{arg\,max}_{\Psi\in\mathcal{P}\left(\textsc{Prompt}(\mathcal{D})\right)}p(\mathcal{D}\mid\mathcal{L}^{\star}(\Psi))p(\mathcal{L}^{\star}(\Psi))p(\Psi)$
where $\mathcal{P}\left(\cdot\right)$ is the powerset , $p(\Psi)$ is a prior favoring fewer predicates, and we have made explicit the dependence of $\mathcal{L}^{\star}$ upon the predicates $\Psi$ (Section 5.2). But the above objective is intractable for two reasons, addressed as follows (Section A.5): Expensive outer loop: The powerset is exponentially large. Rather than exhaustively enumerate every subset of predicates, we do a local search starting from $\Psi=\varnothing$ and greedily adding new predicates from the LLM. Expensive inner loop: Scoring a candidate subset of predicates is expensive, requiring variational inference and causal process structure learning (Sections 5.1 and 5.2). We therefore run structure learning and parameter estimation only once, using all proposed predicates, and cache the resulting processes and parameters for reuse when scoring different subsets of predicates.
@@ 6 Experiments
Experimental Setup. We evaluate eight approaches across five simulated robotics environments, illustrated in Figure 4, using the PyBullet physics engine (Coumans & Bai, 2016). All results are averaged across three random seeds. For each seed, we train the agent with one or two training tasks, each of which includes one demonstration. We then evaluate their performance on 50 held-out test tasks, which include more objects and more complex goals. In each online learning iteration, the agent performs 8 rollouts in a training task with each rollout lasting a maximum of 300 timesteps.
@@
Figure 7: Performance metrics for various agents across different domains.
@@
(Q1). Our approach consistently outperforms the VLM planning (ViLa), HRL (MAPLE) and STRIPS-style operator learning and planning (VisPred) approaches, achieving a near-perfect score across all domains. For each environment, ExoPredicator learns 1-4 exogenous processes and converges after at most three online interaction iterations (fig. 7(b)). Once trained, it can solve nearly all tasks in Coffee, Domino, and Fan, and over 80% in Boil and Grow. In comparison, we find that MAPLE is unable to achieve a high level of success even with 1000 times more interactions and evaluated only on the training distribution. We hypothesize that this is largely due to the challenge of exploration with sparse rewards. ViLa (zero-shot) was not able to do well in any domains. The few-shot variant achieves good performance in simple domains where satisficing plans share significant similarity with the demonstration plan (e.g., in Coffee, one simply needs to perform Pour and NoOp more times than in the demonstration). We observe that its performance degrades significantly in other domains where it must identify additional rules through trial and error (e.g., the requirement of matching colors in Grow) or domains that require compositional generalization (sequencing skills in potentially new ways). VisPred also struggles in these tasks because it learns in a highly constrained model space. It attempted to learn the exogenous processes as different operators for the NoOp skill, but fell short due to an overly strong inductive bias. This bias restricts preconditions to only include atoms with variables already present in an operator’s effects or option, which is especially limiting for the NoOp skill, where “robot” is the only variable. Moreover, it does not learn about the varying delays for different processes, a feature crucial for effective and efficient planning (e.g., in Boil). We note that ExoPredicator is unable to solve all tasks in Boil, partly because it failed to recognize the full condition under which a spill can happen: it learned a process for water spilling when there is nothing under the faucet, but not when the jug underneath is full.

### v3

@@ Abstract
Long-horizon embodied planning is challenging because the world does not only change through an agent’s actions: exogenous processes (e.g., water heating, dominoes cascading) unfold concurrently with the agent’s actions. We propose a framework for abstract world models that jointly learns (i) symbolic state representations and (ii) causal processes for both endogenous actions and exogenous mechanisms. Each causal process models the time course of a stochastic cause-effect relation. We learn these world models from limited data via variational Bayesian inference combined with LLM proposals. Across five simulated tabletop robotics environments, the learned models enable fast planning that generalizes to held-out tasks with more objects and more complex goals, outperforming a range of baselines.
@@ 1 Introduction
The standard planning representation used for decades, PDDL, also fails to capture this: classical PDDL treats actions as instantaneous; while PDDL 2.1(Fox & Long, 2003) adds durative actions, it still does not explicitly model autonomous exogenous processes without extensions such as PDDL+(Fox & Long, 2002). Learning a symbolic PDDL planning model (Silver et al., 2023; Liang et al., 2024) combinatorially explodes with respect to temporal granularity. Vision-language models (VLMs) and vision-language-action models (Team et al., 2023; Black et al., 2025) could in principle reason about external causal mechanisms, but generalize poorly to novel situations, particularly when reasoning about temporal physical constraints.
@@ 3 Abstracting States, Time, and Causal Processes
Thus an abstract state is the collection of symbolic facts the agent believes about the world. This style of state abstraction is standard in robot task planning (Garrett et al., 2021). We write $\mathcal{S}$ for the set of possible abstract states.
Causal Processes: Informal intuition. A causal process coarsely models a cause-effect relation. For example, opening a faucet above a bucket triggers a chain of cause-effect relations: a stopper opens, hidden pipes fill with water, water rises in the bucket, and eventually the bucket is filled. For decision making, many such details are irrelevant, or omitted from the state abstraction. A causal process therefore abstracts away such details by saying that certain conditions (the “cause”) later lead to other conditions (the “effect”). We further distinguish two kinds of causal processes. Endogenous processes correspond to the agent’s high-level actions or skills. They represent operations that are under the agent’s direct control, such as switching a faucet on. Exogenous processes describe the background dynamics of the environment. They represent processes that unfold autonomously once certain conditions hold, independent of whether the conditions were created by the agent or by some external mechanism, such as a kettle filling with water after being placed under a running tap. This separation allows the agent to reason about the consequences of its actions while also anticipating changes initiated by the environment itself. Figure 2 (top right) shows concrete examples of such causal processes (e.g., switching a burner on, filling a jug), with their conditions, delayed effects, and whether they are endogenous or exogenous. [^1] Our terminology differs from standard definitions in causality (Pearl, 2009) and recent RL literature (e.g., Efroni et al. (2021)), where “exogenous” typically refers to variables that are determined outside the system or are irrelevant to the task. In contrast, our exogenous processes are task-relevant and can be triggered by the agent (and are thus causally downstream), but differ in that they evolve autonomously without requiring continuous agent actuation. This aligns with the definition of “processes” in temporal planning frameworks like PDDL+ (Fox & Long, 2002)
@@
Abstract world models are equipped with a set of causal processes $\mathcal{L}$. Each causal process $L\in\mathcal{L}$ is defined by a schema tuple, $\langle\textsc{Par},C,O,E,W,p^{\text{delay}}\rangle$ (see fig. 2, top right, for concrete schema examples), where:
  • Par (parameters) is a list of typed variables present in the condition or effect of the process.
@@
  • $W\in\mathbb{R}$ (log strength) quantifies how likely the effect will happen when the conditions are satisfied.
  • $p^{\text{delay}}$ is a probability distribution over the delay between the process’s activation and effect (Section C.3).
Endogenous processes further include a skill $c\in\mathcal{C}$. A schema is instantiated into a ground causal process $\underline{L}=\langle\underline{\textsc{Par}},\underline{C},\underline{O},\underline{E},W,p^{\text{delay}}\rangle$ (and optionally $\underline{c}$) by substituting its parameters with specific objects. For example, the endogenous process SwitchBurnerOn has parameters ?robot and ?burner. Its condition $C$ requires that the burner is off and the robot’s hand is free. Once executed, its effect $E$ (the burner being on) occurs after a delay sampled from $p^{\text{delay}}$.
Interdependence of state abstraction and causal processes.
The right abstractions depend on downstream tasks, as our goal is generalization to unseen planning problems. Tasks demanding a more detailed state abstraction will generally require modeling more temporal dynamics. This couples the causal processes to the state abstraction.
Probabilistic Semantics.
The causal processes define how abstract states evolve over time. Mathematically, $\mathcal{L}$ defines a probabilistic generative model over sequences of abstract states $\left\{s_{t}\right\}$ at fine-grained timesteps $t\in\mathbb{N}$. At each step, active causal processes contribute pressure for their effect atoms to change, while a separate frame term prefers all atoms to remain as they are. To write this compactly, we use indicator functions for the conditions at start: $C_{\underline{L}}(s_{1:t})$ is $1$ iff all atoms in the condition-at-start set $C$ of ground process $\underline{L}$ are satisfied at state $s_{t}$ but not satisfied at state $s_{t-1}$, and 0 otherwise; similarly, $O_{L}(s_{t^{\prime}+1:t-1})$ is $1$ iff all atoms in the overall-condition set $O$ hold at every step between $t^{\prime}+1$ and t1t-1. We probabilistically model the effect of ground process $\underline{L}$ as a potential $E^{j}_{\underline{L}}$ upon each feature $j$ of the abstract state:
@@ 5 Process Learning and Predicate Invention
We initialize with 1-2 demonstration trajectories, and then perform online learning (Figure 3). Online learning involves planning to solve training tasks to collect further trajectories (when planning fails, random actions are taken). At each stage of online learning, we have a dataset of state-action trajectories $\mathcal{D}_{\text{low}}=\{(x_{0},\underline{c}_{0},\ldots,\underline{c}_{T-1},x_{T})\}$. Given predicates $\Psi$, this generates a dataset of abstract state-action trajectories $\mathcal{D}_{\text{abs}}=\{(s_{0},\underline{c}_{0},\ldots,\underline{c}_{T-1},s_{T})\}$. Given this dataset, we learn predicates (Section 5.3), exogenous processes (section 5.2), and continuous parameters (Section 5.1)—described in a reverse order, because each component is used in the previous component during learning.
@@ 5.1 Parameter Learning via Variational Inference
Each causal process has continuous parameters for the delay distribution $p^{\text{delay}}_{L}$ and the probabilistic weight $W_{L}$, and we have a global parameter $W_{F}$ for the frame axiom. Fixing all discrete structure—including the predicates—we seek parameters maximizing the marginal likelihood $p(\mathcal{D}_{\text{abs}})=p(s_{1:t}|s_{0})=\sum_{\left\{\Delta_{1:t}^{\underline{L}}\right\}}p(s_{1:t},\{\Delta_{1:t}^{\underline{L}}\}|s_{0})$. This is intractable because there are combinatorially many ways of timing the cause-effects.
We therefore approximate the marginal likelihood by introducing variational distributions over the time at which each process realizes its effect: categorical distribution $q_{t}^{\underline{L}}(A_{t}^{\underline{L}})$ encodes belief about $A^{\underline{L}}_{t}$, the “arrival time” of the effect coming from the process $\underline{L}$ due to its activation at time $t$. This is a change of basis from delays to absolute times, $A^{\underline{L}}_{t}=\Delta_{t}^{\underline{L}}+t$, which leads to a tidier ELBO decomposition. We provide the full expression and derivation of the following variational lower bound on the marginal likelihood in Section A.3, which we optimize using Adam (Kingma, 2014).
$\log p(s_{1:t})\geq\underbrace{\mathbb{E}_{q}\!\left[\log p\big(\{A_{t}^{\underline{L}}\}\big)\right]}_{\text{delay model}}\;+\;\underbrace{\mathbb{E}_{q}\!\left[\log p\big(\{s_{t}\}\mid\{A_{t}^{\underline{L}}\}\big)\right]}_{\text{abstract dynamics}}\;+\;\underbrace{\mathbb{H}_{q}\!\left[\{A_{t}^{\underline{L}}\}\right]}_{\text{entropy regularizer}}$
@@ 5.2 Learning Exogenous Processes: Bayesian Model Selection, LLM Guidance
Given a set of trajectories $\mathcal{D_{\text{abs}}}$, we would ideally learn the causal processes $\mathcal{L}^{\star}$ maximizing
$\mathcal{L}^{\star}$ $=\operatorname*{arg\,max}_{\mathcal{L}}p(\mathcal{L}|\mathcal{D_{\text{abs}}})=\operatorname*{arg\,max}_{\mathcal{L}}p(\mathcal{D_{\text{abs}}}|\mathcal{L})p(\mathcal{L})$ (intractable)
where $p(\mathcal{L})$ is a minimum description length prior and $p(\mathcal{D}_{\text{abs}}|\mathcal{L})$ is approximated by Section 5.1. As this optimization is intractable, we learn a separate process $L_{\mathcal{C}}$ for each cluster $\mathcal{C}$:
$\mathcal{L}^{\star}$ $=\bigcup_{\mathcal{C}\in\textsc{Cluster}(\mathcal{D}_{\text{abs}})}\{L_{\mathcal{C}}\}\quad\text{where }L_{\mathcal{C}}=\operatorname*{arg\,max}_{L}p(\mathcal{C}|L)p(L)$ (still intractable)
@@
$\mathcal{L}^{\star}$ $=\bigcup_{\mathcal{C}\in\textsc{Cluster}(\mathcal{D}_{\text{abs}})}\left\{L_{\mathcal{C}}\right\}\quad\text{ where }L_{C}=\operatorname*{arg\,max}_{L\in\textsc{Prompt}(\mathcal{C})}p(\mathcal{C}|L)p(L)$ (tractable)
@@ 5.3 Learning State Abstractions: Program Synthesis and Local Search
$\Psi^{\star}$ $=\operatorname*{arg\,max}_{\Psi\in\mathcal{P}\left(\textsc{Prompt}(\mathcal{D}_{\text{low}})\right)}p(\mathcal{D}_{\text{low}}\mid\mathcal{L}^{\star}(\Psi))p(\mathcal{L}^{\star}(\Psi))p(\Psi)$
where $\mathcal{P}\left(\cdot\right)$ is the powerset, $p(\Psi)$ is a prior favoring fewer predicates, and we have made explicit the dependence of $\mathcal{L}^{\star}$ upon the predicates $\Psi$ (Section 5.2). But the above objective is intractable for two reasons, which we address as follows (Section A.5):
  • Expensive outer loop: The powerset is exponentially large. Rather than exhaustively enumerate every subset of predicates, we do a local hill-climbing search starting from $\Psi=\varnothing$ and greedily adding new predicates from the LLM (Silver et al., 2023).
  • Expensive inner loop: Scoring a candidate subset of predicates is expensive, requiring variational inference and causal process structure learning (Sections 5.1 and 5.2). We therefore run structure learning and parameter estimation only once, using all proposed predicates, and cache the resulting processes and parameters for reuse when scoring different subsets of predicates.
@@ 6 Experiments
Experimental Setup. We evaluate eight approaches across five simulated robotics environments, illustrated in Figure 4, using the PyBullet physics engine (Coumans & Bai, 2016). All results are averaged across five random seeds. For each seed, we train the agent with one or two training tasks, each of which includes one demonstration. We then evaluate their performance on 50 held-out test tasks, which include more objects and more complex goals. In each online learning iteration, the agent performs 8 rollouts in a training task with each rollout lasting a maximum of 300 timesteps.
@@
Figure 7: Performance metrics for various agents across different domains. The error bars/shaded regions show $\pm$ 1 standard deviation.
@@
(Q1). Our approach consistently outperforms the VLM planning (ViLa), HRL (MAPLE) and STRIPS-style operator learning and planning (VisPred) approaches, achieving a near-perfect score across all domains. For each environment, ExoPredicator learns 1-4 exogenous processes and converges after at most three online interaction iterations (fig. 7(b)). Once trained, it can solve nearly all tasks in Coffee, Domino, and Fan, and over 80% in Boil and Grow. In comparison, we find that MAPLE is unable to achieve a high level of success even with 1000 times more interactions and evaluated only on the training distribution. We hypothesize that this is largely due to the challenge of exploration with sparse rewards. ViLa (zero-shot) was not able to do well in any domains. The few-shot variant achieves good performance in simple domains where satisficing plans share significant similarity with the demonstration plan (e.g., in Coffee, one simply needs to perform Pour and NoOp more times than in the demonstration). We observe that its performance degrades significantly in other domains where it must identify additional rules through trial and error (e.g., the requirement of matching colors in Grow) or domains that require compositional generalization (sequencing skills in potentially new ways). VisPred also struggles in these tasks because it learns in a highly constrained model space. It attempted to learn the exogenous processes as different operators for the NoOp skill, but fell short due to an overly strong inductive bias. This bias restricts preconditions to only include atoms with variables already present in an operator’s effects or option, which is especially limiting for the NoOp skill, where “robot” is the only variable. Moreover, it does not learn about the varying delays for different processes, a feature crucial for effective and efficient planning (e.g., in Boil). We note that ExoPredicator is unable to solve all tasks in Boil, partly because it failed to recognize the full disjunctive condition under which a spill can happen: it learned a process for water spilling when there is nothing under the faucet, but not when the jug underneath is full.